OpenStudy (anonymous):

5/13t=-9 how do i do did step by step?

OpenStudy (er.mohd.amir):

cross multiply first

OpenStudy (er.mohd.amir):

that is 5=-9*13t

OpenStudy (er.mohd.amir):

-9*13t=5

OpenStudy (er.mohd.amir):

-13t=5/9

OpenStudy (er.mohd.amir):

-t=5/(9*13)

OpenStudy (er.mohd.amir):

t=-5/117

OpenStudy (anonymous):

thank you